OpenStudy (anonymous):

WILL MEDAL AND FAN YOU PLESE I NEED TO GET THIS DONE BY MIDNIGHT 3sqrt(12)xsqrt(6)

OpenStudy (anonymous):

it's multiplying it not a variable for the x

jimthompson5910 (jim_thompson5910):

I'm using the rule \(\Large \sqrt{xy} = \sqrt{x}*\sqrt{y}\) \[\Large 3*\sqrt{12}*\sqrt{6}\] \[\Large 3*\sqrt{12*6}\] \[\Large 3*\sqrt{72}\] \[\Large 3*\sqrt{36*2}\] \[\Large 3*\sqrt{36}*\sqrt{2}\] \[\Large 3*6*\sqrt{2}\] \[\Large 18*\sqrt{2}\] ------------------------------------------------------- So,\[\Large 3*\sqrt{12}*\sqrt{6}=18*\sqrt{2}\]

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOU!! can you help me with like 5 more please?

jimthompson5910 (jim_thompson5910):

I'll help with 2 more

OpenStudy (anonymous):

ok let me pick which 2 i need the most help on.

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

ok 4/sqrt5

OpenStudy (anonymous):

\[\frac{ 4 }{ \sqrt{5} }\]

jimthompson5910 (jim_thompson5910):

Multiply top and bottom by \(\large \sqrt{5}\) to rationalize the denominator \[\large \frac{4}{\sqrt{5}}\] \[\large \frac{4*\sqrt{5}}{\sqrt{5}*\sqrt{5}}\] \[\large \frac{4\sqrt{5}}{\sqrt{5*5}}\] \[\large \frac{4\sqrt{5}}{\sqrt{25}}\] \[\large \frac{4\sqrt{5}}{5}\] ------------------------------------------------------- Therefore, \[\large \frac{4}{\sqrt{5}}=\frac{4\sqrt{5}}{5}\]

jimthompson5910 (jim_thompson5910):

Notice how I used the rule \(\Large \sqrt{xy} = \sqrt{x}*\sqrt{y}\) again

OpenStudy (anonymous):

ok i know you said you'd only help with one more after this but can you help two? i got the other two now, but there's still two more im really, really stuck on :(

jimthompson5910 (jim_thompson5910):

just one more

jimthompson5910 (jim_thompson5910):

the idea is to use them to answer the others (I don't want to do all your hw)

OpenStudy (anonymous):

what if you help help with one then just check the other? pleaseeee!!! this is an extra credit i really need an A in this class:(

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

THANK YOU! :) ok this is the one i just need you to check, then we can do the other

OpenStudy (anonymous):

\[3\sqrt{6}-4\sqrt{6}\]

OpenStudy (anonymous):

\[12\sqrt{6}\]

jimthompson5910 (jim_thompson5910):